OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 7 liters of natural oil for every 5 liters of synthetic oil. If 623 liters of natural oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (anonymous):

7 liters of natural oil = 5 liters of synthetic oil 1 liters of natural oil=5/7 liters of synthetic oil 623 liters of natural oil=623*5/7 liters of synthetic oil

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Petrolyn oil = natural oil + synthetic oil

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

445?

OpenStudy (anonymous):

Petrolyn oil = 623 liters of natural oil + 623*5/7 liters of synthetic oil

OpenStudy (anonymous):

that dont make since i did 623 and added it to 5/7

OpenStudy (anonymous):

623 + 623*5/7 = 1068

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

i didnt add 2 623s ok got it

OpenStudy (anonymous):

Great.